OpenStudy (anonymous):

is (infinity/0) an indeterminate quotient? where we can apply L'Hopital's rule?

OpenStudy (rulnick):

yes, quite indeterminate.

OpenStudy (ash2326):

Yeah it's indeterminate quotienent We can apply in form of limits such as \[\infty / \infty and 0/0\]

OpenStudy (rulnick):

L'Hopital is about limits of differentialble functions.

OpenStudy (rulnick):

"Differentialble." Great. I ken spel.

OpenStudy (anonymous):

can you use L'H rule for a infinity/0 problem?

OpenStudy (anonymous):

I thought it had to be 0/0 or infinity/infinity for L'H rule to apply

OpenStudy (rulnick):

No, I don't think so, it wouldn't help except to confirm an undefined limit.

OpenStudy (zarkon):

you cant use L'hopitals rule on those...and anywhay ... why would you want to

OpenStudy (anonymous):

heres my thing...lim as x-->0 of (cotx-1)/x

OpenStudy (zarkon):

if it is \[\infty/0\] then there are only 3 possibilities either the limit is infinity,-infinity, or does not exits

OpenStudy (anonymous):

The answer is suppose to be 0, but I got infinity lol. let me try again

OpenStudy (anonymous):

thanks guys!

OpenStudy (zarkon):

it is not 0

OpenStudy (zarkon):

it is infinity

OpenStudy (anonymous):

it was originally lim as x-->0 (cotx -1/x) it was a limit of differences so i had to change it up

OpenStudy (anonymous):

i think i changed it wrongly?

OpenStudy (zarkon):

cot(x)-(1/x) as x->0 is zero

OpenStudy (zarkon):

(cot(x)-1)/x is infinity as x-> 0

OpenStudy (anonymous):

wait so i don't have to change it up to apply "L'HR"?

OpenStudy (anonymous):

i changed it to (sinx-1)/(xcosx) which look more like something we could use L'H Rule

OpenStudy (zarkon):

\[\frac{x\cos(x)-\sin(x)}{x\sin(x)}\]

OpenStudy (anonymous):

oh dammit! i mixed up tan with cot...thanks lol!! but wait so if i get something like 0/infinity or infinity/0 that means i changed up the function wrong right? and have to have a better means to change it to apply L'H Rule?

OpenStudy (zarkon):

0/infinity or infinity/0 are possible...but dont require l'hospitals rule \(\frac{0}{0}\) or \(\frac{\pm\infty}{\pm\infty}\) to use the rule

OpenStudy (anonymous):

alright thanks everyone!!!